OpenStudy (anonymous):

Which equation best represents the line? (4 points) y=1/3x-1 y=3x-1 y=-x+1/3 y=3x+1

OpenStudy (godlovesme):

line?

OpenStudy (godlovesme):

can u pick points?

OpenStudy (anonymous):

give me a sec i try lol

OpenStudy (godlovesme):

lol ok

OpenStudy (anonymous):

0,0 -1,-1 -2,-2 -3,-3 i think im not good at this

OpenStudy (godlovesme):

umm not really look at this ones (0,-1) and (1,2)

OpenStudy (anonymous):

ok i found them

OpenStudy (godlovesme):

now u see the point more clearly :D how many ups and overs does it take u to go from one point to another?

OpenStudy (anonymous):

what is up and over ?? what u mean by dat

OpenStudy (godlovesme):

hole a sec ...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3 ups 1 over

OpenStudy (godlovesme):

yessss :) awesome job so that's slope rise/run --> up/over so what's our slope here?

OpenStudy (anonymous):

1/3 ?

OpenStudy (godlovesme):

up is 3 over is 1 slope =rise/run=up/over= 3/1 = 3 :)

OpenStudy (godlovesme):

do u get it? @shadow2239

OpenStudy (anonymous):

D i think

OpenStudy (godlovesme):

close but no what's the y-intercept of the line?

OpenStudy (anonymous):

1 i think

OpenStudy (godlovesme):

-1 that's where the line hits the y-axis do u see it ?

OpenStudy (anonymous):

ya thats what i meant whoops lol

OpenStudy (godlovesme):

;) lol so what's ur answer?

OpenStudy (anonymous):

y =3x-1

OpenStudy (anonymous):

?

OpenStudy (godlovesme):

yep good job gurl :)

OpenStudy (anonymous):

yay thanks

OpenStudy (godlovesme):

u r welcome, as always :D